OpenStudy (anonymous):

sec^2(t)sin(t) in terms of cos(t) in Q2, I'm getting (sqrt(1-cos^2(t)))/(cos^2(t)), does that look right?

sam (.sam.):

Try this

sam (.sam.):

Try using \[\sec^2t=\frac{1}{\cos^2 t}\]

OpenStudy (anonymous):

What getting is that sec^2(t) is, \[\frac{ 1 }{ \cos ^{2}(t) } \]and that sin(t) in Q2 is,\[\sqrt{1-\cos ^{2}(t)}\] So the final answer is by multiply together is \[\frac{ \sqrt{1-\cos ^{2}(t)} }{ \cos ^{2}(t) }\] Is that correct?

OpenStudy (loser66):

yes, I am with you